A baseball player, Mickey, who bats 310 (or .310) gets an average of 3.1 hits in ten at bats. We will assume that each time Mickey bats he has a 0.31 probability of getting a hit. This means Mickeys at bats are independent from one another. If we also assume Mickey bats 5 times during a game and that x= the number of hits that Mickey gets then the following probability mass function, p(x), and cumulative distribution function F(x) are reasonable:

x 0 1 2 3 4 5
p(x)      .156 .351 .316 a .032 .003
F(x)      .156 .507 .823 b .997

1.00

a. Calculate the variance of 6x.
b. Assume Mickey bats exactly 5 times (a 5-at-bat-game) in each of four consecutive games. What is the probability he gets 2 hits in exactly 2 of those 4 games? (Getting exactly 2 hits will be called a 2-hit-game.)
Answer #1

from given data:

x P(X=x) xP(x) x2P(x)
0 0.156 0.00000 0.00000
1 0.351 0.35100 0.35100
2 0.316 0.63200 1.26400
3 0.142 0.42600 1.27800
4 0.032 0.12800 0.51200
5 0.003 0.01500 0.07500
total 1.5520 3.4800
E(x) =μ= ΣxP(x) = 1.5520
E(x2) = Σx2P(x) = 3.4800
Var(x)=σ2 = E(x2)-(E(x))2= 1.0713

a) variance of 6x =Var(6x) =62Var(x) =36*1.0713 =38.5668

b)

P(2 hits) =0.316

therefore from binomial distribution :

probability he gets 2 hits in exactly 2 of those 4 games =P(X=2)=(4C2)*(0.316)2(1-0.316)2 =0.2803